### v1

# Minimal 3-regular Penny Graph

Abstract
We prove that a 3-regular penny graph has at least 16 vertices and show that such a graph with 16 vertices exists.
1 Introduction
The first author met the second author during the All-Soviet Math Olympiad in Yerevan in 1974. Alexander was one year older than Tanya. In 1976, Tanya was still competing as a high school senior, while Alexander was already a freshman at Moscow State University. Alexander proposed the following two related puzzles for the 1976 Moscow Olympiad [2].
Puzzle 1. You are given a finite number of points on a plane. Prove that there exists a point with at most 3 closest neighbors.
Just in case, by closest neighbors we mean all points at the minimal distance from a given point. This puzzle is connected to our main discussion, so we provide its solution in Section 2. We encourage the reader to solve it before reading our solution.
This puzzle implies that if a planar configuration is such that every point has the same number of closest neighbors, then that number is no greater than 3. This is where the second problem comes in.
Puzzle 2. Can you place a finite number of points on the plane in such a way that each point has exactly $3$ closest neighbors?
The second problem has an elegant solution with $24$ points chosen from a triangular grid, as seen in Figure 1.
Figure 1: A solution with 24 points
By the way, Tanya competed in the 1976 Moscow Olympiad and was given the first problem to solve; she is sure she solved it. The story continued almost $40$ years later, when Alexander emailed Tanya a configuration with $16$ points in Figure 2 and the following conjecture.
Figure 2: A solution with 16 points
Conjecture 1 (Karabegov’s Conjecture).
Any finite planar point configuration in which every point has exactly $3$ closest neighbors must contain at least $16$ points.
In this paper, we prove the conjecture. We start with preliminary observations and definitions.
2 Preliminaries
We consider a finite number of points on a plane.
As we promised, we provide a solution to the first puzzle.
We prove it by contradiction. Suppose we built a set $M$ of points on a plane, so that each point has at least $4$ closest neighbors. Suppose $d$ is the smallest among all these distances. Suppose $M^{\prime}\subset M$ is the subset of all points such that the distance from these points to their closest neighbors is $d$. Obviously, the set $M^{\prime}$ is non-empty, and each point in it has at least $4$ closest neighbors, all of which must belong to $M^{\prime}$.
Now, we build the convex hull of $M^{\prime}$, which is a convex polygon. We denote it by $H$. Suppose $A$ is one of the extreme points in $H$. To remind, an extreme point on a convex hull is a point that is not a convex linear combination of any other points of the hull. The point $A$ is a vertex of $H$ and the internal angle at $A$ is less than $180^{\circ}$.
Let $B_{1}$, $B_{2}$, $B_{3}$, and $B_{4}$ be the $4$ points in $M^{\prime}$ that are at distance $d$ from $A$. Consider two distinct neighbors $B_{i}$ and $B_{j}$ of $A$. Any angle $\angle B_{i}AB_{j}$ has to be at least $60^{\circ}$, as $|B_{i}B_{j}|\geq d$. This contradicts our conclusion that $B_{1}$, $B_{2}$, $B_{3}$, and $B_{4}$ all fit inside an interior angle of less than $180^{\circ}$.
We are now back to building a set in which each point has exactly 3 closest neighbors. We consider all such sets. Suppose we found such a set with the minimal total number of points. We show that there can be at most one distance that realizes the closest neighbors.
Proposition 2.
The closest distance is unique within the minimal set.
Proof.
Suppose we have several such distances. Pick the minimum one, which we denote by $d$. Pick the subset $S$ of points, such that they have the closest neighbors at a distance $d$. Each point in $S$ will have 3 closest, all of which belong to $S$. Thus, we can remove all other points and get a smaller configuration $S$. ∎
We are now in a situation where the closest distance is the same for every point. We can assume this distance to be equal to one. We consider our points and connect the ones at distance $1$ with an edge, thus obtaining a graph.
Our graph is a penny graph. A penny graph is a graph where two vertices are connected by an edge if and only if their distance is the minimum distance among all pairs of vertices. Such graphs are called penny graphs because we can place a penny at each vertex, so that the coins will not overlap and will touch each other if and only if the corresponding vertices are connected. It is assumed that the diameter of a penny is $1$. There is a large body of research on penny graphs (see, e.g., [1, 3, 4], and references therein).
It is clear that a penny graph is a unit distance graph, which is defined as a graph formed from a collection of points in the Euclidean plane by connecting two points whenever the distance between them is exactly $1$.
Moreover, a penny graph is a matchstick graph: a graph that can be drawn in the plane in such a way that its edges are line segments with length $1$ that do not cross each other. That is, it is a graph that has an embedding that is simultaneously a unit distance graph and a planar graph. Figure 3 shows two unit distance graphs: the one on the right is also a matchstick graph.
Figure 3: Two unit distance graphs with a matchstick graph on the right
The graph on the left, the unit distance graph, has a planar embedding, but it does not have an isometric planar embedding, so it is not a matchstick graph.
Matchstick graphs allow two vertices to be at a smaller distance than the length of a matchstick. Thus, not every matchstick graph can be a penny graph, as, in a penny graph, the distance between any two vertices not connected by edges is longer than the length of an edge.
In our graph, each vertex has exactly $3$ neighbors. Such graphs are called $3$-regular graphs.
By the degree sum formula, a $3$-regular graph with $n$ vertices has $\frac{3n}{2}$ edges. In particular, the number of vertices $n$ must be an even number.
To conclude, we want to prove that a 3-regular penny graph with 16 vertices, shown in Figure 2, is minimal.
3 Building Blocks
In our construction, we will use two unit distance graphs as building blocks. We call them a triangle and a rhombus. Figure 4 shows both.
Figure 4: Triangle and rhombus
Both constructions in Figures 1 and 2 consist of rhombuses connected to each other. The $24$-point construction consists of $6$ rhombuses, while the $16$-point construction consists of $4$ rhombuses.
What will happen if we try a similar construction with $3$ rhombuses? The image in Figure 5 shows such a configuration, which now has extra edges with the shortest distance. We now see $3$ points with more than three closest neighbors each, violating the condition.
Figure 5: A construction with 33 rhombuses
4 Boundary
A planar graph divides the plane into regions. We define the boundary of a graph as the set of its edges that are incident to the outer region. We call the edges on the boundary the boundary edges.
Let the boundary of our 3-regular penny graph be an $n$-gon whose sides are edges of length 1. In Figure 6 we give an example of a portion of the boundary made of thick edges, where the edges not incident to the boundary are not shown. Each boundary edge is adjacent to two non-boundary edges. We say that the sum of the angles between a boundary edge and the adjacent non-boundary edges is the angle contribution of the boundary edge. The sum of angle contributions of all the boundary edges is equal to the sum of the interior angles of the boundary $n$-gon, which is $180^{\circ}(n-2)$.
Figure 6: Different types of boundary edges and vertices
Suppose that there are $k$ sides of this boundary polygon that are edges of equilateral triangle faces of the graph. We call such boundary edges the triangle edges, and we call the other boundary edges free edges. Each triangle edge contributes $120^{\circ}$ to the sum of interior angles. Now we want to look at the angle contribution of the free edges.
But before that, we introduce a geometric lemma.
Lemma 3.
Let $KABL$ be a quadrilateral such that $|KA|=|AB|=|BL|=1$, $\angle KAB>60^{\circ}$, $\angle ABL>60^{\circ}$, and $\angle KAB+\angle ABL<180^{\circ}$. Then $|KL|<1$.
We asked ChatGPT to prove this lemma and verified the proof. The proof is technical and not very elegant, so we omit it. Instead, we show an image illustrating Lemma 3 in Figure 7, so the statement is easier to follow.
LLBBAAKK
Figure 7: The quadrilateral in Lemma 3
Lemma 4.
A free edge contributes at least $180^{\circ}$ to the sum of the interior angles.
Proof.
Let $AB$ be a free edge. Each vertex on the boundary has exactly one non-boundary edge coming out of it. Let $AK$ and $BL$ be such non-boundary edges. The fact that $AB$ is a free edge guarantees that the vertices $K$ and $L$ are distinct. Therefore $|AL|>1$ and $|BK|>1$, which imples that $\angle ABL>60^{\circ}$ and $\angle KAB>60^{\circ}$. We also know that $KABL$ is a quadrilateral with $|KA|=|AB|=|BL|=1$. The free edge $AB$ contributes the angle $\angle KAB+\angle ABL$ to the sum of the interior angles. Suppose that this contribution is less than $180^{\circ}$. Then the quadrilateral $KABL$ satisfies the assumptions of Lemma 3, implying that $|KL|<1$. This contradicts the fact that no two vertices are at distances shorter than $1$. ∎
We are ready for our main proof.
Theorem 5.
A minimal 3-regular penny graph has $16$ vertices.
Proof.
We proved that each triangle edge contributes $120^{\circ}$ to the sum of edges of the boundary polygon. Lemma 4 shows that a free edge contributes at least $180^{\circ}$.
Recall that our boundary polygon has $n$ edges with $k$ triangle edges. By subtracting the angle contribution of the triangle edges, we get that the total angle contribution of the nkn-k free edges is equal to 180(n2)120k180(n-2)-120k degrees. On the other hand, it is at least 180(nk)180(n-k) degrees. Therefore, we have the inequality
$180(n-2)-120k\geq 180(n-k),$ (1)
which simplifies to
$k\geq 6.$
Assume that $r$ boundary triangles are parts of rhombuses. Since each boundary rhombus has $2$ boundary edges and $4$ vertices, we can say that each one of its boundary edges contributes $2$ vertices to the graph. Each boundary triangle that is not a part of a rhombus contributes $3$ vertices. Hence, all boundary triangles will contribute 2r+3(kr)=3kr2r+3(k-r)=3k-r vertices.
Suppose the number of vertices is less than $16$, and, using the fact that the number of vertices is even, we thus suppose that the total number of vertices is at most $14$. Then $3k-r\leq 14$, implying that $r\geq 3k-14$. Now adding the fact that $k\geq 6$, we get that $r\geq 4$. Therefore, there are at least two boundary rhombuses.
We will show that any free edge adjacent to a rhombus contributes more than $180^{\circ}$ to the sum of interior angles. Suppose this edge is $AB$, with $AK$ and $BL$ non-boundary edges. Suppose that the vertex $B$ belongs to a rhombus, implying that $L$ is also part of that rhombus and has $3$ edges inside the rhombus. If $\angle KAB+\angle ABL=180^{\circ}$, then $KABL$ is a parallelogram, and $|KL|=1$, implying that $L$ has an extra edge, leading to a contradiction, as shown in Figure 8.
KKLLAABB
Figure 8: The boundary in Theorem 5
Since there are free boundary edges adjacent to boundary rhombuses, the inequality (1) is strict and, therefore, $k>6$, which implies that $r>4$ and thus there are at least $3$ boundary rhombuses which contribute $12$ vertices to the graph.
We now show that our boundary cannot consist of 3 rhombuses and nothing else.
Indeed, if we have $3$ boundary rhombuses and each boundary vertex belongs to a rhombus, then there are $3$ free edges connecting these rhombuses. The total angle contribution of these free edges will be exactly $540^{\circ}$, and therefore each free edge will contribute $180^{\circ}$. As we showed before, this leads to extra edges coming out of the internal vertices of the rhombuses as seen in Figure 5. Thus, the graph is not 3-regular.
It follows that we need to add at least one more point to the boundary. If we add a point to the boundary, we need to add more points to the inside. And one extra point on the inside needs to have at least three edges connecting it to new vertices. Thus, the number of vertices is at least $16$. ∎
Figure 2 shows a $3$-regular penny graph with $16$ vertices. Thus, $16$ is the smallest number of vertices in a $3$-regular penny graph.
5 Acknowledgments
When we started writing this paper, we did not know the names of unit distance, matchstick, and penny graphs. We also did not know whether our theorem was new. We asked ChatGPT, but it was not helpful. After we figured out the names, we found several experts in penny graphs and asked whether our theorem was known; they said no. We are grateful to David Eppstein, Arsenii Sagdeev, and Konrad Swanepoel for their prompt replies.